# MAC: A Benchmark for Multiple Attribute Compositional Zero-Shot Learning

### v2

@@ Abstract
Compositional Zero-Shot Learning (CZSL) aims to learn semantic primitives (attributes and objects) from seen compositions and recognize unseen attribute-object compositions. Existing CZSL datasets focus on single attributes, neglecting the fact that objects naturally exhibit multiple interrelated attributes. Their narrow attribute scope and single attribute labeling introduce annotation biases, misleading the learning of attributes and causing inaccurate evaluation. To address these issues, we introduce the Multi-Attribute Composition (MAC) dataset, encompassing 22,838 images and 17,627 compositions with comprehensive and representative attribute annotations. MAC shows complex relationship between attributes and objects, with each attribute type linked to an average of 82.2 object types, and each object type associated with 31.4 attribute types. Based on MAC, we propose multi-attribute compositional zero-shot learning that requires deeper semantic understanding and advanced attribute associations, establishing a more realistic and challenging benchmark for CZSL. We also propose Multi-attribute Visual-Primitive Integrator (MVP-Integrator), a robust baseline for multi-attribute CZSL, which disentangles semantic primitives and performs effective visual-primitive association. Experimental results demonstrate that MVP-Integrator significantly outperforms existing CZSL methods on MAC with improved inference efficiency.
1 Introduction
Figure 1: Comparison of previous dataset and ours. MAC provides comprehensive and representative attributes for the objects.
@@ 1 Introduction
UT-Zappos [50] 16 12 149 29,126 24.75 24.75 136$\times$102
MIT-States [15] 115 245 1,962 53,753 8.01 17.06 256$\times$256
C-GQA [30] 413 674 7,767 39,298 18.81 11.52 286$\times$223
MAC (Ours) 99 259 17,627 22,838 31.41 82.18 1092$\times$982
Table 1: Comparison of existing CSZL datasets. “Avg.attr” represents the average number of attribute types per object type, while “Avg.obj” denotes the average number of object types per attribute type.
Compositional zero-shot learning (CZSL) [1, 22, 26, 2, 18] aims to compose knowledge of learned semantic primitives (e.g., attributes and objects) to recognize unseen compositions during the testing phase, where each attribute and object has been encountered during training. It highlights the model’s capacity to generalize knowledge and infer novel concepts from known primitives, mimicking human cognitive abilities. Additionally, it serves as the foundation for various downstream tasks such as open-vocabulary object recognition [9, 6, 28, 4, 45, 36, 43] and human object interaction [48, 8, 46, 56, 44].
Existing CZSL datasets, such as MIT-States [15], UT-Zappos [50], and C-GQA [30], provide single-attribute annotations, neglecting the fact that real-world objects typically exhibit multiple interrelated attributes. This single-attribute annotation significantly hinders the model’s understanding of the object’s attributes, as single-label classification losses often suppress correct but unlabeled attributes. Single-attribute annotation also leads to inaccurate evaluation of the model performance. It not only causes valid predictions to be misjudged as incorrect but also hinders a comprehensive evaluation of model performance. Even when a prediction does not match the single-attribute ground truth, it may still be valid. For example, the object in Fig. 1 can be described as “yellow orange” based on its color or as “fresh juicy orange” based on its state. Besides, these datasets neglect the co-occurrence relationship of attributes. For instance, attributes like “cool” and “refreshing” frequently appear together in images of ice-cold cola. Introducing such correlations could enhance model representation learning. Additionally, the images of these datasets for the same object category often demonstrate similar appearance features resulting in low complexity in composition space as shown in Tab. 1.
@@
To address these limitations, we design a new dataset: Multi-Attribute Composition (MAC) dataset. MAC includes 22,838 images and 17,627 compositions, with 60.76% of the images annotated with three or more attributes. Compared to existing datasets, MAC is better suited for CZSL due to following characteristics. (1) Comprehensive Annotation. Comprehensive attribute annotation helps the model learn attribute more effectively and reduces misjudgments during evaluation. However, identifying all suitable attributes from a given set is challenging. We initialize three attribute candidate sets (search keywords, CLIP, and GPT-4o) and ensure completeness and accuracy through manual checking and manual supplementation. (2) Rich Image Diversity. Existing datasets primarily feature common attribute-object compositions, neglecting sparse relationships. For example, while concepts like “blue strawberry” are scarcely observed in real-world scenarios, such unusual compositions can provide hard samples for evaluating compositional reasoning capabilities. Current datasets also exhibit visual homogeneity within each object category, limiting models’ ability to learn robust attribute associations. We further analyze this phenomenon in section 1.2 of the Appendix. MAC utilizes LLM to generate diverse compositions for each object type, then crawls images from search engines to ensure visual variety. (3) Representative and Generalizable Attributes. The attribute set directly determines the scope of the dataset. Attributes must be representative to effectively describe an object’s appearance. Additionally, they should generalize across diverse objects to prevent models from inferring attributes solely based on object categories. MAC leverages LLM knowledge to carefully select 99 representative and generalizable attributes, spanning low-level patterns (e.g., “grainy”) to functional states (e.g., “folded”). As in Tab. 1, each attribute type in MAC can be applied to 82.18 object types on average, showing complex attribute-object relationship, significantly outperforming existing datasets.
We propose multiple-attribute compositional zero-shot learning to better evaluate the model’s compositional understanding. We design 6 metrics to measure the accuracy and comprehensiveness of the model prediction results. To ensure consistency with existing works, we report results on AUC, Seen, and Unseen using algorithms from [30]. Nine methods are implemented on MAC, covering graph models [30], prompt learning [58, 31], semantic disentanglement [11, 24] and LVM [23]. Moreover, we propose Multi-attribute Visual-Primitive Integrator (MVP-Integrator), which firstly eliminates composition input and leverages dual-branch prompt tuning to disentangle primitives than applies self-attention [40] for primitive association. Extensive experiments show that MVP-Integrator significantly outperforms existing methods with improved efficiency. In summary, the main contributions of our work are three-fold:
  • We propose multi-attribute compositional zero-shot learning that requires comprehensive understanding of interdependent attributes and objects, extending CZSL to real-world multi-label scenarios.
@@
  • We create the Multi-Attribute Composition (MAC) dataset with detailed attribute annotations and diverse images addressing the limitations of annotation bias and visual homogeneity in existing CZSL datasets. We evaluate several representative methods on it.
  • We propose MVP-Integrator which firstly eliminates composition input. Experimental results show that it achieves SOTA performance on MAC with improved efficiency.
@@
(b) Binding relations.
(c) Attributes co-occurrence.
@@
Figure 2: Dataset Statistics. Our dataset covers images in comprehensive and diverse attributes with rich associations. (a) shows the proportion of images with different numbers of attribute labels; (b) illustrates the binding relationships between attributes and objects. The top section displays the distribution of attributes across varying numbers of associated objects, while the bottom section depicts the reverse distribution; (c) shows the co-occurrence of the top 15 attributes (inner circle) with their most frequently associated attributes (outer circle) ; (d) displays the number of images per primitive (attribute and object) for MAC.
@@
2 Related Work
Compositional Zero-Shot Learning.
@@ 2 Related Work
Compositional Zero-Shot Learning (CZSL) [37, 14, 38, 18, 39, 20] aims to recognize unseen attribute-object compositions at test time while each semantic primitive exists in the training sample. Existing methods for CZSL without large-scale pre-trained models can be divided into two main categories. One group of methods train two separate classifiers to predict attribute and object respectively [22, 53, 42, 54, 17, 16, 13]. Another group of methods learns the joint attribute-object semantic representation of seen and unseen components using transformation functions in a shared feature space [1, 29, 30, 41]. However, learning semantic primitives from scratch is difficult because of the complexity of the compositions. The recent work is starting to research how to transfer pre-trained vision-language models to composition prediction tasks [51, 7, 34, 57, 19, 49, 25, 3, 10, 21]. CSP [31] first adapts CLIP [34] by replacing the text vocabulary with trainable tokens. DFSP [24] proposes a cross-modal decomposed fusion module to learn more expressive image features. Troika [11] adopts a Multi-Path framework to dynamically adjust prompts and performs image-to-text feature fusion. In this work, we propose MVP-Integrator which is the first CZSL method that eliminates the composition branch and utilizes self-attention for visual-primitive association.
CZSL Datasets and Multi-attribute Datasets.
@@
CZSL datasets like MIT-States [15], UT-Zappos [50], and C-GQA [30] are commonly utilized for studying CZSL. UT-Zappos is a relatively simple dataset focusing on the composition of shoes, materials and brands. MIT-States characterizes state variations within image classes using a diverse set of adjectives and their antonyms. C-GQA includes numerous attributes and objects, derived from images cropped from the GQA [12] based on the bounding boxes. The labels of these three datasets contain only one attribute. This misleads the model’s understanding of attributes and results in inaccurate evaluation. Datasets such as COCO Attributes [32], PACO [35], and VAW [33] offer multiple-attribute annotations. However, they do not guarantee comprehensive coverage of relevant attributes, and they exhibit trivial attribute-object relationships, with many objects being associated with a narrow set of attributes, limiting compositional complexity. To tackle these problems, we construct a new dataset by collecting high-quality images that exhibit diverse appearances variation for each object category. We ensure comprehensive attribute annotations through multiple strategies and propose a multiple-attribute compositional learning task.
3 MAC benchmark
3.1 Dataset Overview
@@
The MAC dataset comprises 22,838 images and 17,627 compositions, with 99 attribute types and 259 object types. The MAC dataset is divided into training, validation, and testing sets with 14,591 , 3,007 and 5,240 images respectively. Our images encompass a variety of representative attributes and objects, with each object associated with multiple attributes.
@@
Fig. 2(a) shows the multi-attribute properties of MAC, with 60.76% of the images annotated with three or more attributes. Fig. 2(b) illustrates the complex relationship between attributes and objects in the MAC. The attributes in MAC are not exclusively linked to specific objects; 53% of the attributes can applied to more than 60 types of objects. The bottom part of Fig. 2(b) highlights a clear peak in the mid-range and reveals that 79% of object categories associated with more 25 attributes, demonstrating the rich appearance variations of images within the same object in MAC. Fig. 2(c) illustrates the co-occurrence of different attributes. Frequently co-occurring attributes often share high semantic relevance, such as “cooked” and “viscous” or “smooth” and “silky” . Further analysis of the phenomenon is provided in section 1.1 of the Appendix. Fig. 2(d) demonstrates number of instances per attribute and object. In the MAC dataset, 72.6% of object categories have images within the range of mean $\pm$ std, indicating a relatively balanced distribution.
3.2 Dataset Construction
@@
In the MAC dataset, each image contains a single object label and multiple attribute labels. To create a high-quality dataset with diverse images and comprehensive attribute annotations, we follow a three-step process as illustrated in Fig. 4(a). First, we carefully define the semantic primitives (i.e., attributes and objects) using GPT-4o and manual verification (Primitive Set Definition). We use GPT-4o to generate diverse compositions for each object by selecting suitable attributes. These compositions serve as keywords to retrieve images from search engines, forming the initial dataset (Collect Images of Diverse Compositions). To cover suitable attributes as comprehensive as possible, we use CLIP [34] and GPT-4o to predict attributes for each image. Annotators then select correct attributes for each image based on these predictions and the initial search keywords (Multiple Attributes Labeling). This process results in a high-quality multi-attribute composition dataset with comprehensive attribute labels. Details of the LLM prompts and annotation interface are in Appendix sections 1.3 and 1.4.
Primitive Set Definition. The primitive set defines the object classes and attribute classes, which directly determines the dataset’s scope and significantly influences its quality. Existing datasets, such as MIT-States [15] and C-GQA [30], often include numerous synonyms and non-visual attributes, hindering the learning of semantic primitives and compromising evaluation accuracy. We carefully define the primitive set by leveraging GPT-4o. Take the attribute set as an example, we consider the attribute set of MIT-States as an initial set and refine it through the following steps. First, we ask GPT-4o to eliminate attributes that are not visible or do not directly describe the appearance of the object. Then, we categorize the remaining attributes into mutually exclusive groups and merge synonyms. Finally, we ask GPT-4o to enrich each group with new attributes. To ensure that each attribute has a unique meaning, we merge synonyms using WordNet [27] and perform manual checks to create the primitive set. The object set is constructed in the same manner. The resulting primitive set comprises 99 attributes and 259 objects. We use a refined set to collect diverse images and annotate them in detail.
Collect Images of Diverse Compositions. For a high-quality composition dataset, the diversity of attribute types associated with each object is crucial. It is not enough to simply collect a large number of images for each object type. These images must show the object with different attributes. This diversity is beneficial for learning how attributes and objects interact and preventing the models from inferring attributes solely based on object categories. To achieve this, we create a diverse set of feasible attribute-object compositions. Specifically, using the objects and attributes from the primitive set, we task GPT-4o to generate plausible compositions that include three attributes each. Compared to single attribute compositions, this strategy generates more diverse compositions with attributes from different descriptive perspectives. We also leverage GPT-4o to perform self-evaluation, rating each composition’s logical validity and feasibility on a scale from 1 to 5. Compositions with scores above 3 are selected as the final composition set. We use these compositions as keywords to crawl images from Google. Leveraging the search engine’s robust retrieval capabilities, we are able to collect diverse images with various attributes.
Multiple Attributes Labeling. Multiple attribute labels will provide more useful information for the model. Insufficient attribute annotations will mislead the model into overlooking the correct attributes that actually exist. However, it is significantly hard to discover all suitable attributes from a given set for an image. To ensure comprehensive labeling of attributes, we provide annotators with three reference attribute sets: the keyword set, the GPT-4o set, and the CLIP set. The keyword set consists of the search keywords used to collect images. It contains three attributes plus one object per image, reflecting the data stored in search engines like image tags, file names, text around the images. These terms are relevant and offer a starting point for annotation but may not always accurately represent the image content. For example, the term “frozen” might bring up images of the movie “Frozen” instead of images depicting objects that are literally frozen. The GPT-4o set is generated by GPT-4o, providing more accurate attribute candidates. However, it has hallucinations, occasionally assigning textually plausible but visually incorrect attributes (e.g., predicting a frozen apple as “fresh”). The CLIP set is generated by CLIP, which retrieves attributes for an object $o_{i}$, using prompts {``$a\;photo\;of\;<a_{j}\;o_{i}>^{\prime\prime}\;|a_{j}\in A\}$, where $A$ represents the set of attributes. The top 5 predictions are selected to maximize the coverage of suitable attributes, though some irrelevant attributes may also be included. As shown in Fig. 5, the CLIP set and the GPT-4o set exhibit a distinct complementary relationship. Using both sets ensures more comprehensive attribute annotation. Furthermore, annotators are instructed to supplement other appropriate attributes beyond the three sets and discard images without a prominent object or of poor quality. The proportion of valid attributes in each set and their contribution to the final attribute annotations are presented in Tab. 2.
@@
Candidate Set Valid Proportion Final Contribution
Keyword 12.4% 19.9%
CLIP 14.8% 31.1%
GPT-4o 45.9% 62.8%
Human Add - 20.6%
Table 2: Effectiveness of attribute candidates. “Valid Proportion” indicates the accuracy of attributes in each set, while “Final Contribution” represents their share in the final dataset. The sum of “Final Contribution” exceeds 100% due to overlapping attributes.
3.3 Task Formulation

### v4

@@ Abstract
Compositional Zero-Shot Learning (CZSL) aims to learn semantic primitives (attributes and objects) from seen compositions and recognize unseen attribute-object compositions. Existing CZSL datasets focus on single attributes, neglecting the fact that objects naturally exhibit multiple interrelated attributes. Their narrow attribute scope and single attribute labeling introduce annotation biases, misleading the learning of attribute and causing inaccurate evaluation. To address these issues, we introduce the Multi-Attribute Composition (MAC) dataset, encompassing 22,838 images and 17,627 compositions with comprehensive attribute annotations. MAC shows a complex relationship between attributes and objects, with each attribute type linked to an average of 82.2 object classes, and each object type associated with 31.4 attribute classes. Based on MAC, we propose multi-attribute compositional zero-shot learning that requires deeper semantic understanding and advanced attribute associations, establishing a more realistic and challenging benchmark for CZSL. We propose Multi-attribute Visual-Primitive Integrator (MVP-Integrator), a robust baseline for multi-attribute CZSL, which disentangles semantic primitives and performs effective visual-primitive association. Experiments demonstrate that MVP-Integrator significantly outperforms existing CZSL methods on MAC with improved efficiency. The dataset will be released publicly upon publication. The code and data are available at: https://github.com/xs1317/MAC.
I Introduction
Compositional zero-shot learning (CZSL) [1, 2, 3, 4] aims to compose knowledge of learned semantic primitives (e.g., attributes and objects) to recognize unseen compositions during the testing phase, where each attribute and object has been encountered during training. It highlights the model’s capacity to generalize knowledge and infer novel concepts from known primitives, mimicking human cognitive abilities. Additionally, it serves as the foundation for various downstream tasks such as open-vocabulary object recognition [5, 6] and human object interaction [7, 8].
Existing CZSL datasets, such as MIT-States [9], UT-Zappos [10], and C-GQA [11], provide single-attribute annotations, neglecting the fact that real-world objects typically exhibit multiple interrelated attributes. Single-attribute annotation significantly hinders the model’s understanding of the object’s attributes, as single-label classification losses often suppress correct but unlabeled attributes. Single-attribute annotation also leads to inaccurate evaluation of the model performance. It not only causes valid predictions to be misjudged as incorrect but also hinders a comprehensive evaluation of model performance. Even when a prediction does not match the single-attribute ground truth, it may still be valid. For example, the object in Fig.˜1(a) can be described as “yellow orange” based on its color or as “fresh juicy orange” based on its state. Besides, these datasets neglect the co-occurrence relationship of attributes. For instance, “cool” and “refreshing” frequently appear together in images of ice-cold cola. Introducing such correlations could enhance model representation learning. Additionally, the images of these datasets for the same object class often demonstrate similar appearance features resulting in low complexity in composition space as shown in Tab.˜I.
(a) Comparison to previous datasets.
(b) Task description.
Figure 1: (a) Compared to previous datasets. MAC provides comprehensive and representative attributes for the objects. (b) Multiple attribute compositional zero-shot learning.
TABLE I: Comparison of existing CSZL datasets. “Avg.attr” represents the average number of attribute types per object type, while “Avg.obj” denotes the average number of object types per attribute type.
@@ I Introduction
[…]
MIT-States [9] 115 245 1,962 53,753 8.01 17.06 $\times256$
[…]
@@
To address these limitations, we design a new dataset: Multi-Attribute Composition (MAC) dataset. MAC includes 22,838 images and 17,627 compositions, with 60.76% of the images annotated with three or more attributes. Compared to existing datasets, MAC is better suited for CZSL due to following characteristics. (1) Comprehensive Annotation. Comprehensive attribute annotation helps the model learn attribute more effectively and reduces misjudgments during evaluation. However, identifying all suitable attributes from a given set is challenging. We initialize three attribute candidate sets (search keywords, CLIP, and GPT-4o) and ensure completeness and accuracy through manual checking and manual supplementation. (2) Rich Image Diversity. Existing datasets primarily feature common attribute-object compositions, neglecting sparse relationships. For example, while concepts like “blue strawberry” are scarcely observed in real-world scenarios, such unusual compositions can provide hard samples for evaluating compositional reasoning capabilities. Current datasets also exhibit visual homogeneity within each object class, limiting models’ ability to learn robust attribute associations. We further analyze this phenomenon in section C of the Appendix. MAC utilizes LLM to generate diverse compositions for each object type, then crawls images from search engines to ensure visual variety. (3) Representative and Generalizable Attributes. The attribute set directly determines the scope of the dataset. Attributes must be representative to effectively describe an object’s appearance. Additionally, they should generalize across diverse objects to prevent models from inferring attributes solely based on object classes. MAC leverages LLM knowledge to carefully select 99 representative and generalizable attributes, spanning low-level patterns (e.g., “grainy”) to functional states (e.g., “folded”). As in Tab.˜I, each attribute type in MAC can be applied to 82.18 object classes on average, showing complex attribute-object relationship, significantly outperforming existing datasets.
We propose multiple attribute CZSL to better evaluate the model’s compositional understanding. We use six metrics to evaluate the accuracy and completeness of models in closed-world and open-world settings. For consistency with existing works, we also report results on AUC, Seen, and Unseen following [11]. Nine methods are implemented on MAC, covering graph models [11], prompt learning [12, 13], semantic disentanglement [14, 15] and VLM [16].
Existing CLIP-based methods are suboptimal for multiple attribute CZSL especially in the open-world setting. Their reliance on composition branch leads to a significant performance drop in the open-world setting. The composition branch also increases the inference complexity to $|A|\times|O|$ as it needs to calculate all the text representations for each composition. Finally, these methods often overlook the inherent semantic relationships between attributes, such as the co-occurrence of "shiny" and "clean". To solve these problems, we propose the Multi-attribute Visual-Primitive Integrator (MVP-Integrator), which eliminates composition branch and leverages dual-branch prompt tuning to learn primitives. MVP-Integrator processes the image representations and text representations of primitives jointly. It uses self-attention [17] on concatenated image-text representations to effectively capture attribute-object, attribute-attribute, and image-text relationships. It overcomes the relationship modeling challenges faced by previous methods in the large composition space. Due to the elimination of composition branch, MVP-Integrator inherently achieves consistent performance on open-world and closed-world settings and significantly reduces the inference complexity to $|A|+|O|$. Extensive experiments show that MVP-Integrator outperforms existing methods with improved efficiency. In summary, our main contributions are threefold:
  • We introduce Multi-Attribute CZSL, a new task that extends CZSL to more realistic multi-label scenarios.
@@
  • We construct the MAC dataset, featuring comprehensive attribute annotations and diverse visual content, addressing the annotation bias and visual homogeneity issues in existing CZSL datasets.
  • We propose MVP-Integrator, a novel and efficient CLIP-based approach that removes the composition branch while and achieves state-of-the-art performance.
II Related Work
@@
(b) Attributes co-occurrence.
(c) Binding relations.
@@
Figure 2: Dataset statistics. (a) shows the proportion of images with different numbers of attribute labels; (b) shows the co-occurrence of the top 15 attributes with their most frequently associated attributes; (c) illustrates the binding relationships between attributes and objects. The top section displays the distribution of attributes across varying numbers of associated objects, while the bottom section presents the reverse; (d) displays the number of images per primitive for MAC.
@@
II-A Compositional Zero-Shot Learning.
@@
Compositional Zero-Shot Learning (CZSL) [18, 19, 20, 21] aims to recognize unseen attribute-object compositions at test time while each semantic primitive exists in the training sample. In contrast, conventional ZSL or GZSL methods [22, 23, 24] typically focus on recognizing entirely unseen object classes. While these methods model relationships between local attribute features and object semantics, CZSL poses a more fine-grained scenario that requires understanding complex interactions between attributes and objects. Existing methods for CZSL without large-scale pre-trained models can be divided into two main categories. One group of methods train two separate classifiers to predict attribute and object respectively [25, 26, 27, 28, 29]. Another group of methods learns the joint attribute-object semantic representation of seen and unseen compositions using transformation functions in a shared feature space [1, 30, 31, 32]. However, learning semantic primitives from scratch is difficult because of the complexity of the compositions. The recent work is starting to research how to transfer pre-trained vision-language models to composition prediction tasks [13, 33, 15, 14, 21, 21, 34, 35, 36]. CSP [13] first adapts CLIP [37] by replacing the text vocabulary with trainable tokens. PromptCompVL [33] creates a completely learnable soft prompt including the prefix, state, and object. DFSP [15] proposes a cross-modal decomposed fusion module to learn more expressive image features. Troika [14] adopts a Multi-Path framework to dynamically adjust prompts and performs image-to-text feature fusion. PLO [21] solves composition predicting in a simple to complex manner by utilizing the knowledge of large language models. GIPCOL [34] employs a graph convolutional network to enhance CLIP’s word embeddings, creating text representations optimized for compositional learning. In this work, we propose MVP-Integrator which eliminates the composition branch and utilizes self-attention for visual-primitive association. This design echoes the core idea of mutual-assistance learning as discussed in [38], where task components are encouraged to interact and support each other for improved performance.
II-B CZSL Datasets and Multi-attribute Datasets.
@@
CZSL datasets like MIT-States [9], UT-Zappos [10], and C-GQA [11] are commonly utilized for studying CZSL. UT-Zappos is a relatively simple dataset focusing on the composition of shoes, materials and brands. MIT-States characterizes state variations within image classes using a diverse set of adjectives and their antonyms. C-GQA includes numerous attributes and objects, derived from images cropped from the GQA [39] based on the bounding boxes. The labels of these three datasets contain only one attribute. This misleads the model’s understanding of attributes and results in inaccurate evaluation. Datasets such as COCO Attributes [40], PACO [41], and VAW [42] offer multiple-attribute annotations. However, they do not guarantee comprehensive coverage of relevant attributes, and they exhibit trivial attribute-object relationships, with many objects being associated with a narrow set of attributes, limiting compositional complexity. For example, our analysis of the large-scale VAW dataset, which contains 620 attribute classes and 2,260 object classes, reveals a sparse attribute-object mapping. We observe that over 40% of its attributes are linked to fewer than 20 object classes, and 43.7% of its object classes associated with five or fewer attribute classes. To tackle these problems, we construct a new dataset, MAC, a new dataset constructed specifically to compositional learning. In contart to VAW, the proportions of sparse connections in MAC drop to just 8% and 0%. We achieved this diversity by collecting high-quality images with high visual variation for each object class. We also ensure comprehensive attribute annotations through multiple strategies and propose a multiple attribute compositional learning task.
III Multiple Attribute CZSL Benchmark
III-A Dataset Overview
@@
The MAC dataset comprises 22,838 images and 17,627 compositions, with 99 attribute classes and 259 object classes. The MAC dataset is divided into training, validation, and testing sets with 14,591 , 3,007 and 5,240 images respectively. Our images encompass a variety of representative attributes and objects, with each object associated with multiple attributes.
@@
Fig.˜2(a) shows the multi-attribute properties of MAC, with 60.76% of the images annotated with three or more attributes. Fig.˜2(b) illustrates the co-occurrence of different attributes. Frequently co-occurring attributes often share high semantic relevance, such as “cooked” and “viscous” or “smooth” and “silky” . Further analysis of the phenomenon is provided in section B of Appendix . Fig.˜2(c) illustrates the complex relationship between attributes and objects in the MAC. The attributes in MAC are not exclusively linked to specific objects; 53% of the attributes can applied to more than 60 classes of objects. The bottom part of Fig.˜2(c) highlights a clear peak in the mid-range and reveals that 79% of object classes associated with more 25 attributes, demonstrating the rich appearance variations of images within the same object. Fig.˜2(d) demonstrates number of instances per attribute and object. For object, 76% of classes have image counts within one standard deviation of the mean, indicating a balanced dataset.
III-B Dataset Construction
@@
TABLE II: Effectiveness of attribute candidates. “Attr Num” means the total attribute number of each candidate set. “Valid Proportion” indicates the accuracy of attributes in each set, while “Final Contribution” represents their share in the final dataset. The sum of “Final Contribution” exceeds 100% due to overlapping attributes.
Candidate Set Attr Num Valid Proportion Final Contribution
Keyword 68,514 12.4% 19.9%
CLIP 114,190 14.8% 31.1%
GPT-4o 76,033 45.9% 62.8%
Human Add 11,541 100.0% 20.6%
@@
In the MAC dataset, each image contains a single object label and multiple-attribute labels. To create a high-quality dataset with diverse images and comprehensive attribute annotations, we follow a three-step process as illustrated in Fig.˜4(a). First, we carefully define the semantic primitives (i.e., attributes and objects) using GPT-4o and manual verification (Primitive Set Definition). We use GPT-4o to generate diverse compositions for each object by selecting suitable attributes. These compositions serve as keywords to retrieve images from search engines, forming the initial dataset (Collect Images of Diverse Compositions). To cover suitable attributes as comprehensive as possible, we use CLIP [37] and GPT-4o to predict attributes for each image. Annotators then select correct attributes for each image based on these predictions and the initial search keywords (Multiple Attributes Labeling). This process results in a high-quality multi-attribute composition dataset with comprehensive attribute labels.
Primitive Set Definition. The primitive set defines the object classes and attribute classes, which directly determines the dataset’s scope and significantly influences its quality. Existing datasets, such as MIT-States [9] and C-GQA [11], often include numerous synonyms and non-visual attributes, hindering the learning of semantic primitives and compromising evaluation accuracy. We carefully define the primitive set by leveraging GPT-4o. Take the attribute set as an example, we consider the attribute set of MIT-States as an initial set and refine it through following steps. First, we ask GPT-4o to eliminate attributes that are not visible. Then, we categorize the remaining attributes into mutually exclusive groups and merge synonyms. Finally, we ask GPT-4o to enrich each group with new attributes. To ensure that each attribute has a unique meaning, we merge synonyms using WordNet [43] and manually check the final primitive set. The object set is constructed in the same manner. The resulting primitive set contains 99 attributes and 259 objects.
Collect Images of Diverse Compositions. For a high-quality composition dataset, the diversity of attributes associated with each object is crucial. It is not enough to simply collect a large number of images for each object type. These images must show the object with different attributes. This diversity is beneficial for learning how attributes and objects interact and preventing the models from inferring attributes solely based on object classes. To achieve this, we create a diverse set of feasible attribute-object compositions. Specifically, we use GPT-4o to generate plausible compositions that include three attributes each. Compared to single attribute compositions, this strategy generates more diverse compositions with attributes from different descriptive perspectives. We also leverage GPT-4o to perform self-evaluation, rating each composition’s logical validity and feasibility on a scale from 1 to 5. Compositions with scores above 3 are selected as the final composition set. We use these compositions as keywords to crawl images from Google. Leveraging the search engine’s robust retrieval capabilities, we are able to collect diverse images.
Multiple Attributes Labeling. To address the challenge of comprehensive multi-attribute annotation, our labeling process provides annotators with three reference sets: a keyword set (from image collection), a GPT-4o set, and a CLIP set. The keyword set consists of search keywords used for image collection, containing three attributes and one object per image. These terms reflect data from search engines (like tags and file names), providing a relevant starting point for annotation. However, they may not always accurately represent the image content, as illustrated by the term "frozen" potentially retrieving images from the movie instead of literally frozen objects. The GPT-4o set is generated by GPT-4o, providing more accurate attribute candidates. However, it has hallucinations, occasionally assigning textually plausible but visually incorrect attributes (e.g., predicting a frozen apple as “fresh”). The CLIP set is obtained by using CLIP to retrieve the top-5 attributes for a given image and its object category. As shown in Fig.˜5, the CLIP set and the GPT-4o set exhibit a distinct complementary relationship. Using both sets ensures more comprehensive attribute annotation. Furthermore, annotators are instructed to supplement other appropriate attributes beyond the three sets and discard images without a prominent object or of poor quality. Tab.˜II shows the proportion of valid attributes in each set and their contribution to the final annotations.
III-C Annotation Quality and Potential Data Leakage
Figure 6: Examples of annotation errors. Selected from MVP-Integrator’s top-1 prediction failures, which tend to reveal underlying annotation errors. $\rightarrow$” indicates an object correction. “ADD” denotes a missing attribute. “DELETE” denotes an incorrect attribute.
Annotation Quality. To assess the quality of our annotations, we conducted an error analysis on carefully selected samples. Representative examples of error cases are shown in Fig.˜6. We used our trained MVP-Integrator model to find 200 instances where the top-1 predictions didn’t match the annotations. These instances tended to deviate from the main distribution and are more likely to reflect annotation issues. Our inspection revealed only 7 annotation errors, which we classified into three types: Object Confusion (4 cases), referring to cases where the object was mislabeled as a visually similar or ambiguous one; Missing Attributes (2 cases), where valid attributes were present in the image but not annotated; and Incorrect Attributes (1 case), where the annotated attributes did not align with the actual visual content. The majority of these errors were due to object confusion. Considering that these 200 samples represent the most error-prone subset of the entire dataset, and given the inherent difficulty of comprehensive multi-attribute annotation, this low error rate demonstrates the high quality of our dataset.
Data Leakage Analysis. We’ve also conducted a thorough analysis to address concerns about data leakage from using CLIP in both annotation and training. First, we used diverse strategies for annotations stage to prevent reliance on any single source. As shown in  Tab.˜II, CLIP suggestions made up only 30.9% of the final labels, so our dataset isn’t dominated by its biases. Second, we empirically verify that our model does not disproportionately benefit from CLIP-sourced labels. Only 16.8% of the top-1 correct predictions corresponded to attributes uniquely suggested by CLIP, compared to 30.3% from GPT-4o, showing the model isn’t just memorizing CLIP-derived labels. Finally, we performed an ablation study by retraining our model with all raw CLIP attribute suggestions. The results showed a significant performance decline, with the Exact Match score dropping from 17.42% to 10.81%. This degradation confirms that uncurated CLIP data introduces noise and doesn’t improve learning, further validating that our dataset construction process doesn’t lead to data leakage.
III-D Task Formulation